# Thermodynamics of quantum processes: An operational framework for free energy and reversible athermality

### v1

@@ VI.1 Extractable work from a quantum channel
$\delta W=\delta E-\beta^{-1}\delta S=\delta F_{c}^{\rm eq}.$ (70)
where $F_{c}^{\rm eq}(\rho_{\beta}):=E(\rho_{\beta})-\beta^{-1}S(\rho_{\beta})$ is the equilibrium free energy of the state $\rho_{\beta}$ that is in thermodynamical equilibrium with the bath. An example of such a state is the Gibbs state $\gamma_{\beta}=\frac{e^{-\beta\widehat{H}}}{Z^{\beta}}$. The equilibrium free energy is $F^{\rm eq}_{c}(\gamma_{\beta})=-\beta^{-1}\ln Z^{\beta}$.
@@
$=\delta F^{\beta}_{c}+\beta^{-1}\delta_{i}S.$ (74)
where $F^{\beta}_{c}(\rho)=E(\rho)-\beta^{-1}S(\rho)$ is the non-equilibrium free energy, also called the thermodynamic free energy. Let us now consider the thermalization of a system with Hamiltonian $\widehat{H}$ in the non-equilibrium state $\rho$ to the Gibbs state $\gamma_{\beta}$. This transformation can be carried in two steps:
@@
    $=-F^{\beta}(\rho)-F_{c}^{\rm eq}(\gamma_{\beta})=-F_{c}^{\beta}(\rho).$ (78)
@@
    $\delta W_{\rm rev}$ $=\delta F^{\rm eq}_{c}=F^{\rm eq}_{c}(\gamma_{\beta})-F^{\rm eq}_{c}(\rho)=F^{\rm eq}_{c}(\gamma_{\beta}).$ (79)
@@
$=-F^{\beta}_{c}(\rho)+F_{c}^{\rm eq}(\gamma_{\beta})$ (81)
@@
Partial thermalization: We now consider a bipartite state $\psi_{RA^{\prime}}$ and a quantum channel $\mathcal{N}_{A^{\prime}\to A}$ that acts locally on $A^{\prime}$, say $\rho_{RA^{\prime}}:=\mathcal{N}(\psi_{RA^{\prime}})$. We assume that the channel output and its reference are noninteracting, $\widehat{H}_{RA}=\widehat{H}_{R}\otimes\mathbbm{1}_{A}+\mathbbm{1}_{R}\otimes\widehat{H}_{A}=\widehat{H}_{R}+\widehat{H}_{A}$. We can extract work from this system $\rho_{RA}$ via partial thermalization to the state $\rho_{R}\otimes\gamma^{\beta}_{A}$ (equal to $\psi_{R}\otimes\rho_{A}$), with the following steps (see Fig. 1):
@@
In the following theorem, we establish a direct relation between the resource-theoretic free energy $F^{\beta}[\mathcal{N}]$ of a quantum channel $\mathcal{N}{A^{\prime}\to A}$ and its maximal extractable work via partial thermalization.
@@
The proof follows directly from Proposition 5. The theorem quantitatively establishes the direct relation of the resource-theoretic free energy of a channel with its thermodynamic utility in terms of extractable work.
@@ VII Discussion
Another future direction is to study a resource theory of athermality of quantum channels under other classes of physically and operationally motivated free operations, for example, superchannel generalization of thermal operations [12, 23]. The set of all superchannels that are superchannel generalization of thermal operations will be a subset of Gibbs preserving superchannels. We want to see if the associated resource theory will be reversible or irreversible in the light of [16]. Another future direction would be to inspect the behavior of free energy functions for quantum Markovian and non-Markovian dynamics and their implications on the capabilities of the dynamics to perform thermodynamic tasks [82, 83, 84, 85, 86].
@@ Acknowledgements.
The authors thank Manabendra Nath Bera, Swati Choudhary, Karol Horodecki, and Uttam Singh for discussions. D.G.S. thanks the Department of Science and Technology, Government of India, for the INSPIRE fellowship. S.D. acknowledges support from the Science and Engineering Research Board (now ANRF), Department of Science and Technology, Government of India, under Grant No. SRG/2023/000217 and the Ministry of Electronics and Information Technology (MeitY), Government of India, under Grant No. 4(3)/2024-ITEA. S.D. also acknowledges support from the National Science Centre, Poland, grant Opus 25, 2023/49/B/ST2/02468. S.D. thanks the University of Gdańsk, Gdańsk, Poland for the hospitality during his visit.
@@ A.1 Proof of Theorem 1
$|F[\mathcal{N}]-F[\mathcal{M}]|\leq\frac{1}{\beta}(\varepsilon K+h_{2}(\varepsilon)),$ (90)
@@
$K\geq\max\{D_{\infty}[\mathcal{N}\|\mathcal{T}^{\beta}],D_{\infty}[\mathcal{K}\|\mathcal{T}^{\beta}]\}.$
@@
The condition $\frac{1}{2}\|\mathcal{N}-\mathcal{M}\|_{\diamond}\leq\varepsilon$ implies that $\frac{1}{2}\|\mathcal{N}(\psi_{RA^{\prime}})-\mathcal{M}(\psi_{RA^{\prime}})\|_{1}\leq\varepsilon$ for all $\psi_{RA^{\prime}}$. The continuity of relative entropy in the first argument has been tightened in [87, 88]. Here we use a slightly simpler bound given in [87, Theorem 1], from which it follows that
@@
$K\geq D_{\infty}(\mathcal{N}(\psi_{RA^{\prime}})\|\mathcal{T}^{\beta}(\psi_{RA^{\prime}})).$ (92)
@@
$\lambda_{0}:=\lambda_{\min}(\pi_{R}\otimes\gamma^{\beta}_{A})=\frac{\mathrm{e}^{-\beta E_{\max}}}{|A^{\prime}|Z^{\beta}}.$ (101)
@@
The proof follows from the definition of max-free energy. ∎
@@
which imply $F_{\alpha}^{\beta}[\Theta(\mathcal{N}_{A^{\prime}\to A})]\leq F_{\alpha}^{\beta}[\mathcal{N}_{A^{\prime}\to A}]$. (A2) Reduction to states: For a replacer channel $\mathcal{R}^{\omega}_{A^{\prime}\to A}$, for all $\alpha\in[\frac{1}{2},\infty)$, we have
$\beta F_{\alpha}[\mathcal{R}^{\omega}]$ $=D_{\alpha}[\mathcal{R}^{\omega}\|\mathcal{T}^{\beta}]$
@@
$=D_{\alpha}(\omega_{A}\|\gamma^{\beta}_{A})=\beta F_{\alpha}(\omega).$ (107)
(A3) Minimum value at absolutely thermal channel: This follows from the fact that for $\alpha\in[\frac{1}{2},\infty)$, $D_{\alpha}(\rho\|\sigma)\geq 0$ if both $\rho$ and $\sigma$ are states, and $D_{\alpha}(\rho\|\sigma)=0$ if and only if $\rho=\sigma$. Therefore, given quantum channels $\mathcal{N}_{A^{\prime}\to A}$ and $\mathcal{M}_{A^{\prime}\to A}$, $D_{\alpha}[\mathcal{N}\|\mathcal{M}]=\sup_{\psi\in\operatorname{St}(RA^{\prime})}D_{\alpha}(\mathcal{N}(\psi)\|\mathcal{M}(\psi))=0$ if and only if $\mathcal{N}(\varphi)=\mathcal{M}(\varphi)$ for all $\varphi\in\operatorname{St}(RA^{\prime})$ or $\mathcal{N}=\mathcal{M}$. Therefore, $F_{\alpha}[\mathcal{N}]=\beta^{-1}D_{\alpha}[\mathcal{N}\|\mathcal{T}^{\beta}]\geq 0$ with equality if and only if $\mathcal{N}=\mathcal{T}^{\beta}$.
@@
$\beta F_{\alpha}[\mathcal{N}\otimes\mathcal{M}]=D_{\alpha}[\mathcal{N}\otimes\mathcal{M}\|\mathcal{T}^{\beta}_{1}\otimes\mathcal{T}^{\beta}_{2}].$ (108)
@@
To prove the inequality $\leq$”, we use the following result from [64, Lemma 38 and Proposition 41], for all $\alpha\in[1,\infty)$, for an arbitrary state $\rho_{RA^{\prime}}$, a channel $\mathcal{N}_{A^{\prime}\to A}$ , and a replacer channel $\mathcal{R}_{A^{\prime}\to A}$ which replaces the state with a state $\tau$,
@@
$\leq D_{\alpha}(\mathcal{N}\|\mathcal{R}^{\tau})+D_{\alpha}(\rho_{RA^{\prime}}\|\sigma_{RA^{\prime}}).$ (115)
@@
Similarly, using the quasi-convexity of sandwiched Rényi divergence for $\alpha\in(1,\infty)$, we have the quasi-convexity for free energy,
$F_{\alpha}\left[\sum_{x}p_{x}\mathcal{N}^{x}\right]\leq\max_{x}F_{\alpha}[\mathcal{N}^{x}].$ (122)
@@ A.2 Proof of Lemma 1
For positive semidefinite operators $\rho\neq 0$, $\sigma$ and $\sigma^{\prime}$, if $\sigma^{\prime}\geq\sigma$, we have $D_{\alpha}(\rho\|\sigma^{\prime})\leq D_{\alpha}(\rho\|\sigma)$ for $\alpha\in[\frac{1}{2},\infty)$ ([91, Proposition 4]). Therefore, under the action of a $\mathcal{T}^{\beta}$-subpreserving superchannel, we have
@@ A.3 Proof of Theorem 2
$D_{\alpha}(\sum_{x}p_{x}\psi^{x}_{AB}\|\sum_{x}p_{x}\rho^{x}_{R}\otimes\gamma^{\beta}_{A})\leq\max_{x}D_{\alpha}(\psi^{x}_{RA}\|\rho^{x}_{R}\otimes\gamma^{\beta}_{A}).$ (132)
@@
For the case of $\alpha\to\infty$ Let the equality for the quasi-convexity condition be satisfied, i.e. $D_{\infty}(\rho\|\sigma)=\max_{i}D_{\infty}(\psi_{i}\|\sigma)$, then using Lemma 8
@@ A.4 Proof of Theorem 3
$F^{\varepsilon}_{H}[\mathcal{N}]=\frac{2}{\beta}\mathrm{Dist}^{\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})$ and $F^{\varepsilon}_{\infty}[\mathcal{N}]=\frac{2}{\beta}\mathrm{Cost}^{\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})$.
@@
$=\sup_{m,\Theta}\{\ln m:\penalty 10000\ d_{\Theta_{\psi}^{\Lambda}}((\mathcal{N},\mathcal{T}^{\beta})\rightarrow(\operatorname{id},\mathcal{R}^{\pi}))\leq\varepsilon\}$ (139)
@@
$=\sup_{\psi,\Lambda}\{-\frac{1}{2}\ln\operatorname{tr}(\mathcal{T}^{\beta}(\psi)\Lambda):\penalty 10000\ 1-\operatorname{tr}(\mathcal{N}(\psi)\Lambda)\leq\varepsilon\}$ (141)
$=\sup_{\psi}-\inf_{\Lambda}\{\frac{1}{2}\ln\operatorname{tr}(\mathcal{T}^{\beta}(\psi)\Lambda):\penalty 10000\ 1-\operatorname{tr}(\mathcal{N}(\psi)\Lambda)\leq\varepsilon\}$ (142)
@@
$=\inf_{m}\{\ln m:\penalty 10000\ \frac{1}{2}\|\mathcal{N}-\Theta(\operatorname{id})\|_{\diamond}\leq\varepsilon;\Theta(\mathcal{R}^{\pi})=R^{\gamma}\}$ (145)
$=\inf_{m}\{\ln m:\penalty 10000\ \frac{1}{2}\|\mathcal{N}-\Theta(\operatorname{id})\|_{\diamond}\leq\varepsilon;$
$\hskip 56.9055ptR^{\gamma}=\frac{1}{m}\Theta(\operatorname{id})+\frac{1}{m}\sum_{i}\Theta(\mathcal{W}_{i})\}$ (146)
$=\inf_{\Theta}\{\ln m:\penalty 10000\ \frac{1}{2}\|\mathcal{N}-\Theta(\operatorname{id})\|_{\diamond}\leq\varepsilon;m^{2}R^{\gamma}\geq\Theta(\operatorname{id})\}.$ (147)
@@
$\mathrm{Cost}^{\varepsilon}(\mathcal{N},\mathcal{)}$
$=\inf_{m}\{\ln m:\penalty 10000\ \frac{1}{2}\|\mathcal{N}-\mathcal{E}\|_{\diamond}\leq\varepsilon;m^{2}R^{\gamma}\geq\mathcal{E}\}$ (148)
$=\inf_{m}\{\ln m:\penalty 10000\ \frac{1}{2}\|\mathcal{N}-\mathcal{E}\|_{\diamond}\leq\varepsilon;2\ln m\geq D_{\max}[\mathcal{E}\|\mathcal{T}^{\beta}]\}$ (149)
$=\inf_{m}\{\frac{1}{2}D_{\max}[\mathcal{E}\|\mathcal{T}^{\beta}]:\penalty 10000\ \|\mathcal{N}-\mathcal{E}\|_{\diamond}\leq\varepsilon\}$ (150)

### v2

@@ VI.1 Extractable work from a quantum channel
$\delta W=\delta E-\beta^{-1}\delta S=\delta F_{\rm T}^{\rm eq},$ (70)
where $F_{\rm T}^{\rm eq}(\rho_{\beta}):=E(\rho_{\beta})-\beta^{-1}S(\rho_{\beta})$ is the equilibrium free energy of the state $\rho_{\beta}$ that is in thermodynamical equilibrium with the bath. An example of such a state is the Gibbs state $\gamma_{\beta}=\frac{e^{-\beta\widehat{H}}}{Z^{\beta}}$. The equilibrium free energy is $F^{\rm eq}_{\rm T}(\gamma_{\beta})=-\beta^{-1}\ln Z^{\beta}$.
@@
$=\delta F^{\beta}_{\rm T}+\beta^{-1}\delta_{i}S,$ (74)
where $F^{\beta}_{\rm T}(\rho)=E(\rho)-\beta^{-1}S(\rho)$ is the non-equilibrium free energy, also called the thermal free energy. Let us now consider the thermalization of a system with Hamiltonian $\widehat{H}$ in the non-equilibrium state $\rho$ to the Gibbs state $\gamma_{\beta}$. This transformation can be carried in two steps:
@@
    $=-F^{\beta}(\rho)-F_{\rm T}^{\rm eq}(\gamma_{\beta})=-F_{\rm T}^{\beta}(\rho).$ (78)
@@
    $\delta W_{\rm rev}$ $=\delta F^{\rm eq}_{\rm T}=F^{\rm eq}_{\rm T}(\gamma_{\beta})-F^{\rm eq}_{\rm T}(\rho)=F^{\rm eq}_{\rm T}(\gamma_{\beta}).$ (79)
@@
$=-F^{\beta}_{\rm T}(\rho)+F_{\rm T}^{\rm eq}(\gamma_{\beta})$ (81)
@@
Partial thermalization: We now consider a bipartite state $\psi_{RA^{\prime}}$ and a quantum channel $\mathcal{N}_{A^{\prime}\to A}$ that acts locally on $A^{\prime}$, say $\rho_{RA}:=\mathcal{N}(\psi_{RA^{\prime}})$. We assume that the channel output and its reference are noninteracting, $\widehat{H}_{RA}=\widehat{H}_{R}\otimes\mathbbm{1}_{A}+\mathbbm{1}_{R}\otimes\widehat{H}_{A}=\widehat{H}_{R}+\widehat{H}_{A}$. We can extract work from this system $\rho_{RA}$ via partial thermalization to the state $\rho_{R}\otimes\gamma^{\beta}_{A}$ (equal to $\psi_{R}\otimes\rho_{A}$), with the following steps (see Fig. 1):
@@
In the following theorem, we establish a direct relation between the resource-theoretic free energy $F^{\beta}[\mathcal{N}]$ of a quantum channel $\mathcal{N}_{A^{\prime}\to A}$ and its maximal extractable work via partial thermalization.
@@
The proof follows directly from Proposition 4. The theorem quantitatively establishes the direct relation of the resource-theoretic free energy of a channel with its thermodynamic utility in terms of extractable work [47].
@@ VII Discussion
Another future direction is to study a resource theory of athermality of quantum channels under other classes of physically and operationally motivated free operations, for example, superchannel generalization of thermal operations [12, 23]. The set of all superchannels that are superchannel generalization of thermal operations will be a subset of Gibbs preserving superchannels. We want to see if the associated resource theory will be reversible or irreversible when the uses of catalysis are allowed (cf. [16]). Another future direction would be to inspect the behavior of free energy functions for quantum Markovian and non-Markovian dynamics and their implications on the capabilities of the dynamics to perform thermodynamic tasks [86, 38, 87, 88, 89].
@@ Acknowledgements.
The authors thank Felix C. Binder, Manabendra Nath Bera, Swati Choudhary, Karol Horodecki, Yutong Luo, and Uttam Singh for discussions. D.G.S. thanks the Department of Science and Technology, Government of India, for the INSPIRE fellowship. S.D. acknowledges support from the Science and Engineering Research Board (now ANRF), Department of Science and Technology, Government of India, under Grant No. SRG/2023/000217 and the Ministry of Electronics and Information Technology (MeitY), Government of India, under Grant No. 4(3)/2024-ITEA. S.D. also acknowledges support from the National Science Centre, Poland, grant Opus 25, 2023/49/B/ST2/02468. S.D. thanks the University of Gdańsk, Gdańsk, Poland for the hospitality during his visit.
@@ A.1 Proof of Theorem 1
[…]
@@
$K\geq\max\{D_{\infty}[\mathcal{N}\|\mathcal{T}^{\beta}],D_{\infty}[\mathcal{M}\|\mathcal{T}^{\beta}]\}.$ (91)
@@
The condition $\frac{1}{2}\|\mathcal{N}-\mathcal{M}\|_{\diamond}\leq\varepsilon$ implies that $\frac{1}{2}\|\mathcal{N}(\psi_{RA^{\prime}})-\mathcal{M}(\psi_{RA^{\prime}})\|_{1}\leq\varepsilon$ for all $\psi_{RA^{\prime}}$. The continuity of relative entropy in the first argument has been tightened in [90, 91]. Here we use a slightly simpler bound given in [90, Theorem 1], from which it follows that
@@
$K\geq D_{\infty}(\mathcal{N}(\psi_{RA^{\prime}})\|\mathcal{T}^{\beta}(\psi_{RA^{\prime}}))$ (93)
@@
$\lambda_{0}:=\lambda_{\min}(\pi_{R}\otimes\gamma^{\beta}_{A})=\frac{\mathrm{e}^{-\beta E_{\max}}}{|A^{\prime}|Z^{\beta}},$ (102)
@@
The proof follows from the definition of the max-free energy. ∎
@@
which imply $F_{\alpha}^{\beta}[\Theta(\mathcal{N}_{A^{\prime}\to A})]\leq F_{\alpha}^{\beta}[\mathcal{N}_{A^{\prime}\to A}]$.
(A2) Reduction to states: For a replacer channel $\mathcal{R}^{\omega}_{A^{\prime}\to A}$, for all $\alpha\in[\frac{1}{2},\infty)$, we have
$\beta F^{\beta}_{\alpha}[\mathcal{R}^{\omega}]$ $=D_{\alpha}[\mathcal{R}^{\omega}\|\mathcal{T}^{\beta}]$
@@
$=D_{\alpha}(\omega_{A}\|\gamma^{\beta}_{A})=\beta F^{\beta}_{\alpha}(\omega).$ (108)
(A3) Minimum value at absolutely thermal channel: This follows from the fact that for $\alpha\in[\frac{1}{2},\infty)$, $D_{\alpha}(\rho\|\sigma)\geq 0$ if both $\rho$ and $\sigma$ are states, and $D_{\alpha}(\rho\|\sigma)=0$ if and only if $\rho=\sigma$. Therefore, given quantum channels $\mathcal{N}_{A^{\prime}\to A}$ and $\mathcal{M}_{A^{\prime}\to A}$, $D_{\alpha}[\mathcal{N}\|\mathcal{M}]=\sup_{\psi\in\operatorname{St}(RA^{\prime})}D_{\alpha}(\mathcal{N}(\psi)\|\mathcal{M}(\psi))=0$ if and only if $\mathcal{N}(\varphi)=\mathcal{M}(\varphi)$ for all $\varphi\in\operatorname{St}(RA^{\prime})$ or $\mathcal{N}=\mathcal{M}$. Therefore, $F^{\beta}_{\alpha}[\mathcal{N}]=\beta^{-1}D_{\alpha}[\mathcal{N}\|\mathcal{T}^{\beta}]\geq 0$ with equality if and only if $\mathcal{N}=\mathcal{T}^{\beta}$.
@@
$\beta F^{\beta}_{\alpha}[\mathcal{N}\otimes\mathcal{M}]=D_{\alpha}[\mathcal{N}\otimes\mathcal{M}\|\mathcal{T}^{\beta}_{1}\otimes\mathcal{T}^{\beta}_{2}].$ (109)
@@
To prove the inequality $\leq$”, we use the following result from [67, Lemma 38 and Proposition 41], for all $\alpha\in[1,\infty)$, for an arbitrary state $\rho_{RA^{\prime}}$, a channel $\mathcal{N}_{A^{\prime}\to A}$ , and a replacer channel $\mathcal{R}^{\tau}_{A^{\prime}\to A}$ which replaces all its input states with a state $\tau$,
@@
$\leq D_{\alpha}[\mathcal{N}\|\mathcal{R}^{\tau}]+D_{\alpha}(\rho_{RA^{\prime}}\|\sigma_{RA^{\prime}}).$ (116)
@@
Similarly, for $\alpha\in(1,\infty)$, using the quasi-convexity of sandwiched Rényi divergence, we have the quasi-convexity for the sandwiched Rényi free energy,
$F^{\beta}_{\alpha}\left[\sum_{x}p_{x}\mathcal{N}^{x}\right]\leq\max_{x}F^{\beta}_{\alpha}[\mathcal{N}^{x}].$ (123)
@@ A.2 Proof of Lemma 1
For positive semidefinite operators $\rho\neq 0$, $\sigma$ and $\sigma^{\prime}$, if $\sigma^{\prime}\geq\sigma$, we have $D_{\alpha}(\rho\|\sigma^{\prime})\leq D_{\alpha}(\rho\|\sigma)$ for $\alpha\in[\frac{1}{2},\infty)$ ([94, Proposition 4]). Therefore, under the action of a $\mathcal{T}^{\beta}$-subpreserving superchannel, we have
@@ A.3 Proof of Proposition 2
$D_{\alpha}(\sum_{x}p_{x}\psi^{x}_{AB}\|\sum_{x}p_{x}\rho^{x}_{R}\otimes\gamma^{\beta}_{A})\leq\max_{x}D_{\alpha}(\psi^{x}_{RA}\|\rho^{x}_{R}\otimes\gamma^{\beta}_{A}),$ (133)
@@
For the case of $\alpha\to\infty$, let the equality for the quasi-convexity condition be satisfied, i.e. $D_{\infty}(\rho\|\sigma)=\max_{i}D_{\infty}(\psi_{i}\|\sigma)$, then using Lemma 10,
@@ A.4 Proof of Theorem 2
$F^{\beta,\varepsilon}_{H}[\mathcal{N}]=\frac{2}{\beta}\mathrm{Dist}^{\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})$ and $F^{\beta,\varepsilon}_{\infty}[\mathcal{N}]=\frac{2}{\beta}\mathrm{Cost}^{\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})$.
@@
$=\sup_{\psi,\Lambda}\left\{-\frac{1}{2}\ln\operatorname{tr}(\mathcal{T}^{\beta}(\psi)\Lambda):\penalty 10000\ 1-\operatorname{tr}(\mathcal{N}(\psi)\Lambda)\leq\varepsilon\right\}$ (141)
$=\sup_{\psi}\left[-\inf_{\Lambda}\left\{\frac{1}{2}\ln\operatorname{tr}(\mathcal{T}^{\beta}(\psi)\Lambda):\penalty 10000\ 1-\operatorname{tr}(\mathcal{N}(\psi)\Lambda)\leq\varepsilon\right\}\right]$ (142)
@@
$=\inf_{m}\left\{\ln m:\penalty 10000\ \frac{1}{2}\|\mathcal{N}-\Theta(\operatorname{id})\|_{\diamond}\leq\varepsilon,\penalty 10000\ \Theta(\mathcal{R}^{\pi})=\mathcal{T}^{\beta}\right\}$ (145)
$=\inf_{m}\{\ln m:\penalty 10000\ \frac{1}{2}\|\mathcal{N}-\Theta(\operatorname{id})\|_{\diamond}\leq\varepsilon,$
$\hskip 56.9055pt\mathcal{T}^{\beta}=\frac{1}{m}\Theta(\operatorname{id})+\frac{1}{m}\sum_{i}\Theta(\mathcal{W}_{i})\}$ (146)
$=\inf_{\Theta}\left\{\ln m:\penalty 10000\ \frac{1}{2}\|\mathcal{N}-\Theta(\operatorname{id})\|_{\diamond}\leq\varepsilon,\penalty 10000\ m^{2}\mathcal{T}^{\beta}\geq\Theta(\operatorname{id})\right\}.$ (147)
@@
$\mathrm{Cost}^{\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})$
$=\inf_{m}\left\{\ln m:\penalty 10000\ \frac{1}{2}\|\mathcal{N}-\mathcal{E}\|_{\diamond}\leq\varepsilon,\penalty 10000\ m^{2}\mathcal{T}^{\beta}\geq\mathcal{E}\right\}$ (148)
$=\inf_{m}\left\{\ln m:\penalty 10000\ \frac{1}{2}\|\mathcal{N}-\mathcal{E}\|_{\diamond}\leq\varepsilon,\penalty 10000\ 2\ln m\geq D_{\max}[\mathcal{E}\|\mathcal{T}^{\beta}]\right\}$ (149)
$=\inf_{m}\left\{\frac{1}{2}D_{\max}[\mathcal{E}\|\mathcal{T}^{\beta}]:\penalty 10000\ \|\mathcal{N}-\mathcal{E}\|_{\diamond}\leq\varepsilon\right\}$ (150)